find the equation of the line through (-8,-7) and (0,-2). write the equation in standard form.
Answer by geetha_rama(94)

Point (x1,y1) = (-8,-7) and point (x2,y2) = (0,-2)
Slope of the line m = y2-y1/x2-x1
= (-2 - (-7))/(0-(-8)) = 5/8
Equation of the line is
y -y1 = m(x-x1)
substiuting the value of m and the value of p9oint x1, y1
y - (-7) = (5/8) * (x - (-8))
y+7 = (5/8)*(x+8)
8y+56 = 5x+40
8y = 5x - 16
=>y = (5/8)*x - 2 is the equation of the line